The density of gold is 19300 kg/m3 . A mass of 195 g of gold is shaped into a cube.

What is the length of each edge?

Express your answer using three significant figures. also give the answer in (cm)
Answer #1

Given

The density of the gold is 19300 kg/m3

volume = mass / density

V = (195*10-3 kg) / (19300 kg/m3)

V = 1.01*10-5 m3

volume of the cube V = l3

l = V1/3

L= (1.01*10-5 m3)1/3

l = 0.0216 m = 2.161 cm